OpenStudy (anonymous):

12-4/5(x+15)=4 (PLEASE SHOW WORK)

OpenStudy (anonymous):

Fist distribute: 12 - (4/5)x - 12 = 4 -(4/5)x = 4 x = -5

OpenStudy (anonymous):

how did you get -12 in the first line?

OpenStudy (anonymous):

12-4/5(x+15)=4 12-(4/5)x + (-4/5)(15) = 4 you have to multiply the 15 by -4/5 as well 12 - (4/5)x - 12 = 4

OpenStudy (anonymous):

ahh, thanks.